# Automatic Fact-Checking with Frame-Semantics

### v1

@@ Appendix A Studied Frames
Appendix B Reproducibility
We trained our frame-semantic parser using a NVIDIA GTX 1080 Ti in about 6 hours on a single run without any hyperparameter tuning. Github Copilot was used in the development of some of our code.
B.1 Software License
Our source code and data are released with the MIT License.
@@ B.1 Software License
Table 7: Models and Their Corresponding Hugging Face URLs
B.2 Models
Unless stated otherwise, all models used are the “base” variant. All semantic similarity models are available via the sentence-transformers library. [^9] https://huggingface.co/sentence-transformers Table 7 contains all of the models used and their details.
Claim Predicted Verdict Explanation Fact Check
Table 8: OECD fact check example predictions and explanations
Appendix C Examples
In this section, we provide a few example outputs of our system to give a better understanding of the outputs of our models.
@@ 17 HR 1868 To prevent across-the-board direct spending cuts, and for other purposes Yea Refutes
Table 9: Examples of voting-related claims with the corresponding retrieved bills, votes on the bill, the vote-claim alignment, and fact verification prediction.
C.1 Vote-Claim Alignment
We include two sample claims and their corresponding top three retrieved bills along with the vote-claim alignment for each bill in Table 9.
C.2 Fact Verification Predictions
Vote
We include two claims and their corresponding predictions based on the retrieved bills in Table 9.
OECD
We include two predictions and their explanations on OECD claims in Table 8.
Appendix D LLM Prompts
D.1 Frame-Semantic Parsing
@@
- Side: Additional person(s) involved in the voting on the same Issue alongside the Agent.
- Position: The stance the Agent takes, expressing whether their vote is in favor or against the Issue.
- Frequency: How often the Agent has made this voting decision regarding the Issue.
- Time: When the Agent performed the voting decision.
- Place: Location where the vote took place.
- Support_rate: The percentage of the Agents total votes aligning with the Side over a set period.
@@
- Frame elements should quote exactly from the input
@@
- Dimension: Aspect along which the ranking occurs (e.g., speed, age).
- Comparison_set: Group of entities being compared.
- Time: Time period when the item occupies the rank.
@@
- Dimension: Criterion used for ranking.
- Comparison_set: Set of entities being ranked.
- Time: Time period when the item holds the rank.
@@
- Difference: Amount of change.
- Final_value: Final position on the scale.
- Initial_value: Initial position on the scale.
- Path: Progression between positions on the scale.
- Speed: Rate of change.
- Correlated_variable: Related variable that changes with the Attribute.
- Manner: Method of performing the action.
- Degree: Extent to which the change occurs.
- Circumstances: Context or conditions for the change.
- Result: Outcome of the change.
- Group: Collection of Items undergoing change.
- Time: Time frame during which the change occurs.
- Duration: Length of time for the change to happen.
- Value_range: Range of values along the scale.
- Initial_correlate: State corresponding to the Initial_value.
- Final_correlate: State corresponding to the Final_value.
- Place: Location where the Attribute is measured.
- Initial_state: State of the Item before the change.
- Final_state: State of the Item after the change.
- Period_of_iterations: Duration from start to stop of repeated changes.
- Particular_iteration: Specific instance of a repeated event.
- Containing_event: Broader event encompassing the change.
- Explanation: Rationale for the change.
@@
- Comparison_criterion: Criterion for comparison.
- Degree: Extent of the comparison.
- Time: Period over which the comparison takes place.
@@
- Second_point_in_time: Second time period in comparison.
- Comparison_criterion: Criterion being compared.
- Degree: Extent of the difference.
- Difference: Magnitude of change between the two times.
@@
D.2 Claim-Bill Alignment
@@
D.3 OECD Data Query
@@
- Claim-relevant metrics may appear in columns other than ‘’measure’‘, so reference the claim to identify the appropriate columns to filter by.
@@
- When filtering by ‘’unit_of_measure’‘, ensure to use a standardized metric (e.g., a common currency) to simplify further analysis.
@@
- A list containing cleaned pandas DataFrames based on the claim. If the code is run and the data is not available, return ‘’Data is not Available’‘.
D.4 Fact Verification

### v2

@@ Appendix A Data Quality
Vote Agent, Issue An Agent makes a voting decision on an Issue.
@@
Annotators.
The OECD and voting-related claims were annotated by three students, including one senior PhD student with expertise in frame-semantic parsing. The PhD student provided the final ground truth annotations. The two other students independently provided their annotations, and if they matched the PhD student’s annotations, they were kept. No disagreements occurred during annotation.
@@
Table 7: Models and their corresponding Hugging Face URLs.
Voting claims.
Our dataset of voting-related claims consists of original claims collected from PolitiFact, corresponding URLs, manually cleaned claim text, and references to legislative bills cited in the sources. The dataset contains 79 claims referencing 190 bills, averaging 2.4 bills per claim. However, this distribution is skewed, with a median of only 1 bill per claim due to a few claims referencing a large number of bills.
Fact-checking activity was concentrated in election years, with 86% of the 79 claims fact-checked in 2016, 2018, or 2020. Health insurance was the most frequently mentioned topic, appearing in 11 claims. The terms “for” and “against” were used in 28 and 27 claims, respectively, suggesting a frequent focus on arguments in favor of or opposition to specific policies.
No single individual was disproportionately fact-checked among the 79 claims; Donald Trump had the most fact-checked claims (3), followed by Hillary Clinton (2). Claim lengths ranged from 7 to 62 words, with an average of 18.4 and a median of 17 words.
OECD claims.
The OECD-related claims were more evenly distributed across years, with the top three years (2016, 2014, 2011) accounting for 46% of the 68 claims. The United States was the most frequently mentioned country, appearing explicitly in 25% of the claims. Given PolitiFact’s U.S. focus, implicit references to the U.S. are likely higher.
Taxes were the most common topic, with approximately 20% of the claims mentioning tax-related topics. Claim lengths ranged from 9 to 41 words, with an average of 17.7 and a median of 15 words.
Almost all OECD claims contained a single frame, with only one instance containing multiple frames (Occupy_rank and Comparing_two_entities). The most frequent frames were Comparing_two_entities (25 claims), Occupy_rank_via_superlatives (12 claims), and Occupy_rank (8 claims), indicating a strong emphasis on ranking and comparison within these claims.
@@ 17 HR 1868 To prevent across-the-board direct spending cuts, and for other purposes Yea Refutes
Table 8: Examples of voting-related claims with the corresponding retrieved bills, votes on the bills, vote-claim alignment, and fact verification prediction.
Appendix B Studied Frames
Table 6 provides a comprehensive list of frames studied in this work along with the frame elements used to predict table mappings. The selected frames were chosen based on their relevance to factual claims in the domains of voting and country statistics. Our analysis primarily focused on frames that involve ranking, comparisons, and legislative actions, as these are commonly invoked in fact-checking scenarios.
Appendix C Reproducibility
Ensuring reproducibility is a core principle of this work. We provide details about our experimental setup, computational resources, software dependencies, and models used.
C.1 Computational Resources
We trained our frame-semantic parser on an NVIDIA GTX 1080 Ti GPU. A single training run took approximately 6 hours. No hyperparameter tuning was performed, and the model was trained using default settings unless explicitly stated. The OECD database requires roughly 85GB of storage and the voting records requires 1GB of storage. color=cyan]we don’t train any other models.
C.2 Software and Dependencies
Our implementation is based on PyTorch and Hugging Face’s “transformers” and “sentence-transformers” libraries. A full list of dependencies and their versions is provided in our public repository. GitHub Copilot was used in the development of some of our code to assist with boilerplate functions and standard NLP preprocessing tasks.
C.3 Software License
All source code and dataset annotations are released under the MIT License.
C.4 Models
Unless stated otherwise, all models used in our experiments are the “base” variant. All semantic similarity models are sourced from the “sentence-transformers” library.  [^11] https://huggingface.co/sentence-transformers Table 7 provides an overview of the models used and their parameter count.
Frame-semantic parsing.
Our frame-semantic parsing model is an adaptation of AGED, augmented with the frame classification layer and strategy from Devasier et al. (2024). We fine-tuned this model using a joint loss function to improve frame prediction accuracy. Instead of BERT, as originally used in AGED, we employ RoBERTa for its improved contextual embeddings, which have been shown to enhance performance on frame-based semantic parsing tasks.
Voting agent matching.
Our agent matching function retrieves the BioGuide ID of a political figure from a database. It extracts the agent’s name using frame element spans (FEs) and checks if it is a known presidential nickname, replacing it with the full name if applicable. If not, it uses a SpaCy NER model to extract PERSON named entities. The function first attempts an exact match in the members table, then a fuzzy match using SQL’s LIKE operator. If neither succeeds, it constructs a multi-condition query, incorporating nickname expansions where applicable.
Appendix D Examples
To provide a clearer understanding of our system’s outputs, we include example cases from both the voting and OECD datasets. These examples illustrate how our models process claims, align them with retrieved documents, and generate predictions.
Claim Predicted Verdict Explanation Fact-Check
Table 9: OECD fact-check example predictions and explanations.
D.1 Vote-Claim Alignment
Table 8 presents two sample claims from the voting dataset along with their top three retrieved bills. For each retrieved bill, we include the computed vote-claim alignment score, which indicates the relevance of the bill to the claim. This alignment process plays a crucial role in determining the factual accuracy of voting-related claims.color=red]I don’t any score in the table.
D.2 Fact Verification Predictions
Voting claims.
For voting-related claims, we provide two examples where our model predicts the veracity of claims based on retrieved bills. Table 8 includes the claims, retrieved evidence, and the final prediction score.color=red]no score
OECD claims.
For OECD-related claims, we provide two sample predictions and their corresponding explanations in Table 9. These examples demonstrate how our model interprets numerical comparisons, tax-related claims, and ranking statements to verify factual assertions.
Appendix E LLM Prompts
E.1 Frame-Semantic Parsing
@@
... [omitted for brevity]
@@
- Frame elements should quote exactly from the input.
@@
... [omitted for brevity]
@@
... [omitted for brevity]
@@
... [omitted for brevity]
@@
... [omitted for brevity]
@@
... [omitted for brevity]
@@
E.2 Claim-Bill Alignment
@@
E.3 OECD Data Query
@@
- Claim-relevant metrics may appear in columns other than measure’, so reference the claim to identify the appropriate columns to filter by.
@@
- When filtering by unit_of_measure’, ensure to use a standardized metric (e.g., a common currency) to simplify further analysis.
@@
- A list containing cleaned pandas DataFrames based on the claim. If the code is run and the data is not available, return Data is not Available’.
E.4 Fact Verification